# Identifying heterogeneous micromechanical properties of biological tissues via physics-informed neural networks

### v1

@@ 2.5 Convergence behavior of the PINN training
Fig. 5 shows the convergence histories of the elastic moduli and the training loss of each example. As shown, the training procedures and the estimated elastic moduli converge rapidly within the user-defined duration of $5\times 10^{5}$ epochs. Figs. 5A and D demonstrate the convergence evolution of the training procedure for the first example. The $L^{2}$ relative error of the elastic moduli first decays to below 5% at around 4,000 iterations. The estimated elastic moduli continue to approach the FEA ground truth for the remaining iterations, reaching a final relative error of 0.56%. Similarly, the total training loss decreases from $\mathcal{O}(10^{-1})$ to $\mathcal{O}(10^{-3})$ within the first 150,000 iterations. The PDE and constitutive relation losses have reached a plateau immediately at around 1,000 iterations and consistently maintain $\mathcal{O}(10^{-3})$ and $\mathcal{O}(10^{-4})$ relative error. Similar patterns are observed in the second example (Figs. 5B and E) and C (Figs. 5C and F). The $L^{2}$ relative error of the elastic moduli in the second example decays to below 5% at around 250,000 iterations and steadily reduces the error to 2.91% at the end of the training. The $L^{2}$ relative error of the elastic moduli in example C decays to below 5% at around 150,000 iterations and reduces the error to 2.78% at the end of the training. Across the three examples, total training loss converging behavior aligns more with data loss, suggesting that the training procedure is more sensitive to deviation in the strain reference data.
@@ 3 Discussion
In the current work, we proposed a novel approach to accurately uncover the heterogeneous elastic property of hyperelastic materials using PINNs. We experimented with five FCNN architectures in both methods and identified a top-performing PINN architecture that efficiently produced the best estimations across all examples. We used the same hyperparameters (e.g., loss weights, learning decay rate, and optimization algorithm) in each analysis to ensure generality and that the estimated results are independent of parameter tuning. We have selected network architecture IIB as the top-performing network architecture as it consistently yielded accurate estimations of the full-field elastic modulus, with $L^{2}$ relative errors of $0.43\pm 0.11\%$, $3.28\pm 0.93\%$, and $3.09\pm 0.41\%$, for the three examples respectively, while also took the least computational time.
@@
The positive results of our work have opened new opportunities to accurately quantify the elasticity map in clinical strain-based ultrasound elasticity imaging. Shear wave elastography is a common imaging technique for detecting cancerous tissue by measuring tissue elasticity maps through shear wave velocity. However, commercially available US elastography systems assume that soft tissues have homogeneous and incompressible mechanical properties, despite the high degree of heterogeneity in soft tissue properties [84]. This results in an underestimation of tissue stiffness and less accurate predictions. Our work demonstrated that our architecture can accurately estimate the elastic properties distribution using only strain as the training data. The unique characteristics of PINNs, which allow for the estimation of the local stress and elastic moduli, help circumvent the assumption in the tissue behavior and may provide a more accurate quantitative measure of the stiffness map, therefore, leading to a more accurate diagnosis.
@@
Our method demonstrates promising potential extension to microstructure and material properties studies in a compact, flexible, unified framework. Our network architecture offers immense flexibility to integrate additional empirical data and physical principles relevant to the problem. While the current study focuses on the micromechanical properties of soft materials, our architecture can be easily extended to include data such as the temperature field and thermodynamics conservation laws to study the material microstructure progress during crack propagation, helping accelerate progress in material designs and optimization.
@@
While we have not been focused on optimizing the computational time of our analyses, there are a few approaches one could make to reduce the computational cost. First, our studies adopted the FE nodal coordinates (a total of 5101 nodes) as the PDE training points and reference data collocation points. One may experiment with the effect of reducing the number of training points on estimation accuracy [85, 55, 86]. Second, as noted in Section 2.5, the training loss appears to converge rapidly within the first 250,000 iterations of the analysis; depending on the level of accuracy desired, one may also reduce the number of iterations needed. Third, we used the Adam optimizer in our training; one may experiment with the use of second-order optimizers [87, 64] to achieve faster estimation convergence while maintaining the same order of accuracy. Lastly, there has been active research on methods to reduce GPU memory consumption and wall time in physics-informed machine learning, such as mixed precision [88].
@@
In our work, we have chosen three examples with complex morphology to encapsulate the wide variety of structural patterns in soft tissues. It should be noted that the depth and width of FCNNs may need to be adjusted to accommodate the increasing complexity of the tissue structure. Here, we quantified the structural complexity of the tissue sample using Delentropy. We hope this measure will help guide the designs of the FCNN architecture tailored to individual tissue specimens in future studies.
@@ 4 Conclusions
In conclusion, we have identified a general network architecture that produces highly accurate elastic moduli estimates for complex heterogeneous materials, with $L^{2}$ relative error less than 5% across the examples in our study. The results of this work provide a promising framework for identifying the heterogenous elastic properties in soft materials from images, paving the way to accelerate research progress in material designs, material optimization, and medical applications.
@@ 5 Methods
In this work, we investigated the accuracy of standard PINNs and Fourier-feature PINNs in estimating the elastic properties of heterogeneous materials in 2D. Each PINN method consisted of two variants: one with boundary conditions embedded as hard constraints and one without. Additionally, we established five network architectures with a unique arrangement of independent neural networks. A total of 20 neural network architectures were applied to each example to determine the most effective architecture for predicting the elastic modulus distribution in complex materials. Details on the data generation and details pertaining to the neural network architectures are provided in Sections 5.1 to 5.3.
@@ 5.1 Reference data generation
In tensile tests, the displacement fields often do not provide sufficient local data to uncover the fine features in heterogeneous materials. Therefore, strains ($\epsilon_{ij}$) were used to regulate the data loss and train our neural networks. We generated the strain fields using FEA. The first step was prescribing the heterogeneous structure and elastic properties of a finite element mesh. In the first example, we discretized a unit square domain into 10,000 triangular elements and 5,101 nodes in FEniCS [83]. The heterogeneous pattern was generated by assigning random intensity values to the nodal coordinates using a mean-zero GRF defined as
@@
where $k_{l}(x_{1},x_{2})$ is the Guassian kernal and $l$ is the correlation length. We chose a radial basis kernel function with $l=0.1$ in this work.
In the second and third examples, we utilized the tissue microstructure images obtained from Koos et al. [79] and Weinberg et al. [80] to map the complex microstructure patterns to a finite element mesh. First, we cropped a section of the image and transformed the dimension to a unit square. Then, we used min-max normalization to scale the image intensity values to a range from 0 to 1. We followed the same procedure as the first example, wherein a unit-square geometry was established in FEniCS [83], and discretized the domain into 10,000 cells and 5,101 nodes. Finally, we interpolated the image intensity using cubic interpolation to map the heterogeneous feature of the brain and tricuspid valve leaflet tissue onto the finite element mesh.
In these examples, Young’s modulus of each element was calculated based on the intensity at the nodal coordinates following
@@
where $f(x_{i})$ is the pixel intensity associated with node $i$ on the triangle element; a higher intensity value indicates greater stiffness. We considered a compressible isotropic Neo-Hookean hyperelastic constitutive model [89], where the first Piola-Kirchhoff stress is formulated as

### v3

@@ 2.3.2 Convergence behavior of PINNs
Fig. 5 shows the convergence histories of the elastic moduli and the training loss of each example using network architecture IIB. As shown, the training procedures and the estimated elastic moduli converge rapidly within the user-defined duration of $5\times 10^{5}$ epochs. Figs. 5A and D demonstrate the convergence evolution of the training procedure for the first example. The $L^{2}$ relative error of the elastic moduli first decays to below 5% at around 4,000 iterations. The estimated elastic moduli continue to approach the FEA ground truth for the remaining iterations, reaching a final relative error of 0.56%. Similarly, the total training loss decreases from $\mathcal{O}(10^{-1})$ to $\mathcal{O}(10^{-3})$ within the first 150,000 iterations. The PDE and constitutive relation losses have reached a plateau immediately at around 1,000 iterations and consistently maintain $\mathcal{O}(10^{-3})$ and $\mathcal{O}(10^{-4})$ relative error. Similar patterns are observed in the second example (Figs. 5B and E) and C (Figs. 5C and F). The $L^{2}$ relative error of the elastic moduli in the second example decays to below 5% at around 250,000 iterations and steadily reduces the error to 2.91% at the end of the training. The $L^{2}$ relative error of the elastic moduli in example C decays to below 5% at around 150,000 iterations and reduces the error to 2.78% at the end of the training. Across the three examples, total training loss converging behavior aligns more with data loss, suggesting that the training procedure is more sensitive to deviation in the strain reference data. The estimations and pointwise errors of displacements, strains, and elasticity maps resulting from network IIB are detailed in Appendix B.
@@
2.4 Extended experiments on selected PINN architectures
In Section 2.3, we identified four accurate network architectures that consistently produced accurate estimates of the elasticity map across the test examples. In this section, we first identify the best-performing network architecture by evaluating its estimation accuracy on various heterogeneous tissue patterns. Afterward, we examine the robustness of the selected PINN architecture in the presence of noisy strain data, as well as when extrapolated to different mechanical loading conditions and material constitutive models.
2.4.1 Hetergeneous tissue patterns from ultrasound images
We identified that networks IB, IIB, IC, and IVC consistently yielded accurate full-field elastic parameter prediction of the tissue with less than 5% $L^{2}$ relative error in Section 2.3. Among them, IB and IIB were standard PINNs, while IC and IVC were Fourier-feature PINNs. In the following, we apply these four network architectures to analyze three cases of malignant breast cancer tissues captured from ultrasound images from a public dataset [79]. It is important to note that we used the same hyperparameters as in the previous examples to demonstrate consistency in our methods. However, the hyperparameters could be further refined to achieve better estimation results.
In the three breast cancer tissue examples, we generated the strain data following the same forward FEA procedures as the brain and tricuspid valve tissue examples. In the forward FEA, 40% equibiaxial stretch was applied to the boundary of the specimens. Similarly, the Neohookean hyperelastic material model with Poisson’s ratio of 0.3 was assumed. The full-field elastic parameters were estimated based on the pixel values of the images as described in Section 5.1.
The full-field elastic parameter estimations and pointwise errors of the breast cancer tissues are shown in Fig. 6. These examples have delentropy complexity values ranging from 4.80 to 5.21, which are larger than the values of the previous examples. In cancer tissue example A, the $L^{2}$ relative errors for networks IB, IIB, IC, and IVC were 5.60%, 5.20%, 10.24%, and 3.13%, respectively. For example B, the errors were 12.01%, 11.51%, 13.05%, and 8.48%, respectively. In example C, the errors were 6.92%, 6.69%, 11.80%, and 4.57%, respectively. Across the three examples, network IVC consistently produced accurate prediction with less than 10% $L^{2}$ relative errors, while network IIB consistently delivered the best estimations of full-field elasticity maps within standard PINN architectures. Due to the lower computational time required, network IIB was used in all remaining analyses in the present work.
Figure 6: Breast cancer tissues. Three examples of malignant breast cancer tissues were chosen from a public dataset [79] to demonstrate the applicability of our proposed PINN architectures to a wider range of tissue patterns with delentropy complexity values beyond those in the previous examples. The pointwise absolute errors were all under 0.3 across the three tissue examples and network types. In example (A), the $L^{2}$ relative errors for networks IB, IIB, IC, and IVC were found to be 5.60%, 5.20%, 10.24%, and 3.13%, respectively. Meanwhile, the errors were 12.01%, 11.51%, 13.05%, and 8.48%, respectively in example (B) and 6.92%, 6.69%, 11.80%, and 4.57%, respectively in example (C). Network IVC consistently achieved less than 10% $L^{2}$ relative errors in estimating heterogeneous elasticity maps.
2.4.2 Effect of noise in reference data
Strain data, either from experimental measurements or strain elastography, are typically subject to varying degrees of noise that could interfere with the accuracy of inverse estimation. To assess the robustness of the proposed network architecture, we injected $1\%$, $2\%$, $5\%$, and $10\%$ of white Gaussian noise into the strain tensor of each example. The reference noisy strain data, along with the $L^{2}$ relative errors of the estimated elasticity maps are shown in Fig. 7. Our experimentation revealed that the estimation accuracy of the elasticity map decreased as the level of noise increased in the first example. However, the $L^{2}$ relative errors were similar across different noise levels in the second and third examples. This impartial behavior towards noise measurements could be attributed to the inherently complex structural patterns within the brain and tricuspid valve tissues. Nonetheless, the selected network architecture was able to estimate the elasticity maps with a high degree of accuracy even when the reference data was corrupted with a high level of noise, as the $L^{2}$ relative errors consistently remain below $5\%$ across all examples and noise levels (Fig. 7D).
Figure 7: Effect of noisy reference data. (A-C) The reference strain data of the GRF, brain tissue, and tricuspid valve tissues subjected to $1\%$, $2\%$, $5\%$, and $10\%$ white Gaussian noise. (D) The $L^{2}$ relative errors are below $5\%$ across all examples and noise levels.
2.4.3 Estimating heterogeneous material parameters of plane stress Neo-Hookean model subjected to varying levels of biaxial stretch
In the following experiments, we used the same full-field elasticity maps in Fig. 1 for each example. We modified the material stress-strain behavior to a plane stress Neo-Hookean model and assumed a Poisson’s ratio of 0.45. Each model was subjected to three levels of biaxial stretching ($10\%$, $20\%$, $40\%$) independently, and the reference strain data were obtained using FE simulation. Overall, Network architecture IIB accurately estimated full-field elasticity maps with $L^{2}$ relative errors within $5.04\%$ in all experiments.
Fig. 8 shows the estimated elasticity maps and pointwise errors of the three examples subject to varying levels of biaxial stretch. In the GRF example (Fig. 8A), it was observed that as the percentage of stretch increased, the $L^{2}$ relative errors for the elasticity map estimation also increased. The $L^{2}$ relative errors were found to be $0.17\%$, $0.63\%$, and $0.95\%$ for stretch levels of $10\%$, $20\%$, and $40\%$, respectively. Similar trends of the $L^{2}$ relative errors were observed in the brain tissue example presented in Fig. 8B. The $L^{2}$ relative errors were found to be $2.38\%$, $3.63\%$, and $4.01\%$ for stretch levels of $10\%$, $20\%$, and $40\%$, respectively. However, in the tricuspid valve tissue example (Fig. 8C), higher $L^{2}$ relative errors were found with a lower percentage of biaxial stretch. The $L^{2}$ relative errors were found to be $5.04\%$, $4.25\%$, and $2.23\%$ for stretch levels of $10\%$, $20\%$, and $40\%$, respectively. Moreover, the maximum pointwise error across the estimated full-field elasticity maps generally increased with a higher Delentropy value in the underlying tissue structures. The maximum pointwise error of the full-field elasticity maps across the three examples were 0.03, 0.13, and 0.21, respectively.
Figure 8: Effect of biaxial stretch intensity on PINN predictions and pointwise errors. (A and B) The pointwise error increased as the example underwent larger deformation due to higher prescribed biaxial stretch. On the contrary, (C), higher biaxial stretch produced a more accurate estimate of the full-field elasticity map. In general, the maximum pointwise error of the estimated full-field elasticity maps increased with a higher Delentropy value and complexity in the structural pattern of the examples.
2.4.4 Estimating heterogeneous material parameters of compressible Neo-Hookean, incompressible Mooney Rivlin, and incompressible Gent models
In the following experiments, we modified the material constitutive model to plane stress Neo-Hookean [80], incompressible Mooney Rivlin [81], and incompressible Gent [82]. Detailed formulations of these material models are described in Fig. 5.2. The full-field elasticity maps in Fig. 1 were used to represent the unknown distribution of the material variable $E$, $\mu_{1}$, and $\mu$ in the Neo-Hookean, Mooney Rivlin, and Gent models, respectively. All examples were subjected to $40\%$ biaxial stretch, and the reference strain data were obtained using FE simulation. Network architecture IIB accurately estimated full-field material parameter maps with $L^{2}$ relative errors well within $5\%$ in all experiments, except the tricuspid valve example with Gent model had a $L^{2}$ relative error of $7.05\%$.
Fig. 9 shows the estimated parameter distribution and pointwise errors of the three examples characterized by plane stress compressible Neo-Hookean, incompressible Mooney Rivlin, and incompressible Gent models. In the GRF (Fig. 9A) and tricuspid valve (Fig. 9C) examples, it was observed that the $L^{2}$ relative errors of the heterogeneous material parameter map increased as the complexity of the material constitutive model increased. For the GRF example, the $L^{2}$ relative errors were found to be $0.95\%$, $1.08\%$, and $2.00\%$ for Neo-Hookean, Mooney Rivlin, and Gent, respectively. Similarly, for the tricuspid valve example, the $L^{2}$ relative errors were found to be $2.23\%$, $3.64\%$, and $7.05\%$ for Neo-Hookean, Mooney Rivlin, and Gent, respectively. On the contrary, in the brain tissue example (Fig. 9B), the estimated full-field material parameters of the Gent model were found to have the lowest $L^{2}$ relative error, which were observed to be $4.01\%$, $4.39\%$, and $2.70\%$ for Neo-Hookean, Mooney Rivlin, and Gent, respectively. Consistent with Section 2.4.3, higher Delentropy value in the complexity of the examples yielded a higher pointwise error of the estimated full-field material parameters. The maximum pointwise error of the full-field elasticity maps across the three examples were 0.05, 0.15, and 0.32, respectively.
Figure 9: Effect of material constitutive models on PINN predictions and pointwise errors. (A and C) An increase in the complexity of the mathematical description of the material models correspondingly resulted in an increase in pointwise errors. In both examples, the Gent model reported the highest pointwise errors, while the Neo-Hookean model returned the lowest. Conversely, in (B), the Gent Model had the lowest pointwise errors. Similar to Fig. 8, examples with higher structural complexity produced higher pointwise errors.
@@ 3 Discussion
In the current work, we examined the effects of PINNs for uncovering the heterogeneous material property of nonlinear elastic materials with complex tissue microstructures. We constructed 20 PINN architectures to identify the full-field elastic modulus of soft tissues characterized by an isotropic incompressible Neo-Hookean material model. We used the same hyperparameters (e.g., loss weights, learning decay rate, and optimization algorithm) in each analysis to ensure generality and that the estimated results are independent of parameter tuning. In our computational experiments, we observed that network architecture IIB consistently yielded accurate estimations of the full-field elastic modulus across a wide range of tissue structural patterns and simultaneously took the least computational time. These promising results led to further evaluation of the robustness and generality of architecture IIB on examples subjected to various biaxial stretch intensities and more complex material models, such as the incompressible Mooney Rivlin and Gent models. Across all experimentations, the selected PINN architecture consistently delivered highly accurate estimates of the full-field material constants even in the presence of noisy data. This study presents a critical first step toward a robust PINN model for identifying full-field heterogeneous material constants in soft tissues.
Our proposed approach extends on the works of Kamali et al., and Chen et al.  [53, 54], which focused on identifying the linear elastic properties of heterogeneous materials under small deformation and infinitesimal strains. Here, we applied large deformation to our examples and used three hyperelastic material models to better characterize the material behavior of soft tissues. In addition, our methodology differed from that of Kamali et al. [53] as we utilized full-field strain computations from displacement output to ensure local kinematics were accurately satisfied, instead of using strain fields as network input with boundary tractions to optimize network parameters. Furthermore, our approach eliminates the necessity for stress data, which may not always be readily available and measurable, making it more adaptable to clinical imaging when boundary conditions are not precisely known. While the Poisson’s ratio of soft tissue may vary between 0.3 to 0.5 [83], measuring the compressibility of soft tissue non-invasively is non-trivial. In fact, most clinical strain elastography assumes a constant Poisson’s ratio. To that end, our study focused on identifying the heterogeneous stiffness maps. It is important to note that the current strain data were generated from simulated biaxial numerical experiments, whereas the boundary and loading conditions experienced by tissues in in vivo data could differ depending on the type of examination. Therefore, additional validation of the present framework using medical images will be undertaken.
@@
Our proposed PINN architecture may lead to new non-invasive approaches for quantifying the elasticity map utilizing the full-field strain measurement obtained from magnetic resonance imaging, optical coherence tomography, or strain-based ultrasound elastography. Shear wave elastography is a common imaging technique for detecting cancerous tissue by measuring tissue elasticity maps through shear wave velocity. However, commercially available US elastography systems assume material linearity in soft tissues, despite the high degree of nonlinearity and heterogeneity present [84]. Our work demonstrated that our architecture can accurately estimate the elastic properties distribution using only strain as the training data. PINNs enable simultaneous estimation of the full-field stress and elastic moduli; this characteristic circumvents the need for homogeneous material assumption and may provide a more accurate quantitative measure of the stiffness map, therefore leading to a more accurate diagnosis.
@@
Our method provides a compact, flexible, unified framework for material properties studies. Our network architecture offers immense flexibility to integrate additional empirical data and relevant physical principles and knowledge into the network. For example, the current network could be extended to analyze the material property distributions of anisotropic materials by including additional invariant terms in the material constitutive models. Further, experimental data could also be easily integrated into the network to constrain the optimization path of the network variables, which in turn may further improve the elastic property prediction accuracy. While the current study focuses on the heterogeneous properties of soft materials, our architecture is highly suited for other engineering applications, with an easily extensible framework to include data such as the temperature field and thermodynamics conservation laws to study the material microstructure progress during crack propagation, helping accelerate progress in material designs and optimization.
@@
While optimizing the computational time was not a focus within the scope of this work, there are a few approaches one could make to reduce the computational cost. First, our studies adopted the FE nodal coordinates (a total of 5101 nodes) as the PDE training points and reference data collocation points. One may experiment with the effect of reducing the number of training points on estimation accuracy [85, 49, 86]. Second, as noted in Section 2.3.2, the training loss appears to converge rapidly within the first 250,000 iterations of the analysis; depending on the level of accuracy desired, one may also reduce the number of iterations needed. Third, we used the Adam optimizer in our training; one may experiment with the use of second-order optimizers [87, 57] to achieve faster estimation convergence while maintaining the same order of accuracy. Lastly, there has been active research on methods to reduce GPU memory consumption and wall time in physics-informed machine learning, such as mixed precision [88] which can reduce the computational cost by up to 50%.
In addition to discovering the full-field heterogeneous material parameters, advancement in machine learning has inspired complementary efforts to develop data-driven approaches for identifying material constitutive models appropriate for soft tissues [89]. By discovering constitutive models, we can gain insights into the homogenized, macroscopic material behaviors, providing a mathematical description to best describe phenomenological observation. By discovering full-field material parameters, we can better understand the internal structure and elastic properties of tissues meso- or microscopically, bringing new insights into the behavior of internal tissue structure in response to external mechanical forces that lead to the mechanical responses that were observed macroscopically. Robustly coupling the multiscale mechanical behaviors of soft tissues remains an open challenge [90]. Developing a multiscale machine learning model that couples soft tissue behaviors across macro-, meso-, and microscale to better inform the mechanobiology of soft tissues would be an exciting long-term objective to explore in future studies.
In this context, it is essential to incorporate realistic microstructures derived from high-resolution imaging modalities such as MRI, OCT, or strain elastography. These techniques provide detailed data that can significantly enhance the accuracy and physiological relevance of our models. Integrating this real data will allow us to move beyond randomly distributed material parameters and towards more precise, data-driven representations of tissue microstructure. This approach will not only improve the fidelity of our models but also provide deeper insights into the behavior of biological tissues under various loading conditions, thus better informing the mechanobiology of soft tissues.
@@
Lastly, in our work, we have chosen three examples with complex morphology to encapsulate the wide variety of structural patterns in soft tissues. Here, the structural complexity of the tissue sample was quantified using Delentropy. We hope this measure will help guide the designs of the FCNN architecture tailored to individual tissue specimens. It should be noted that the depth and width of FCNNs may need to be adjusted to accommodate the increasing complexity of the tissue structure. Further, the network output and governing equations can be easily extended to accommodate three-dimensional analysis. However, in the case of 3D analysis, the accuracy of estimation could be compromised if strain measurements are limited to the surface of a specimen. Fortunately, strain values will be accessible throughout the region of interest in strain elastography.
@@ 4 Conclusions
In conclusion, we identified an effective PINN architecture, from 20 architecture variants, that consistently demonstrated excellent estimation accuracy of elasticity maps on three unique heterogeneous materials (GRF, brain tissue, and tricuspid valve tissues). The selected PINN model demonstrated robustness and high estimation accuracy on three different hyperelastic constitutive models (Neo-Hookean, Mooney Rivlin, and Gent), various levels of biaxial stretch, and up to 10% noise in the training data. Further, the selected architecture demonstrated excellent extrapolation to additional soft tissue structures that have Delentropy complexity values beyond the initial experiment examples. The results of this work provide a promising framework for identifying the heterogenous elastic properties in soft materials from images, paving the way to accelerate research progress in material designs, material optimization, and medical applications.
@@ 5 Methods
In this work, we investigated the accuracy of standard PINNs and Fourier-feature PINNs in estimating the material properties of heterogeneous materials in 2D. Each PINN method consisted of two variants: one with boundary conditions embedded as hard constraints and one without. Additionally, we established five network architectures with a unique arrangement of independent neural networks. The current study consists of two parts. First, a total of 20 neural network architectures were applied to each example to determine the most effective architecture for predicting the elastic modulus distribution in complex materials. Second, once a top-performing architecture was identified, it was modified to include noise in the training data and adapted to three different hyperelastic material models, and biaxial stretch levels to assess their performance when extrapolating other mechanical scenarios. Details on the data generation and details pertaining to the neural network architectures are provided in Sections 5.1 to 5.4.
@@ 5.1 Reference data generation
In our experience, the displacement fields often do not provide sufficient local data to uncover the fine features in heterogeneous materials. Therefore, we generated strains ($\epsilon_{ij}$) from FEA to regulate the data loss and train our neural networks. To set up finite element models, we first prescribed heterogeneous patterns and elastic properties to a finite element mesh. In all three examples, the finite element mesh consisted of a unit square with 10,000 equilateral triangular elements and 5,101 nodes.
In the first example, the heterogeneous pattern was generated by assigning random intensity values to the nodal coordinates using a mean-zero GRF defined as
@@
where $k_{l}(x_{1},x_{2})$ is the Guassian kernal and $l$ is the correlation length. We chose a radial basis kernel function with $l=0.1$ in this work. In the second and third examples, we utilized the tissue microstructure images obtained from Koos et al. [71] and Weinberg et al. [72] to map the complex microstructure patterns to a finite element mesh. First, we cropped a section of the image and transformed the dimension to a unit square. Then, we used min-max normalization to scale the image intensity values to a range from 0 to 1. Finally, we interpolated the image intensity using cubic interpolation to map the heterogeneous feature of the brain and tricuspid valve leaflet tissue onto the finite element mesh.
In these examples, Young’s modulus of each element was calculated based on the average nodal intensity value, defined as
@@
where $f(x_{i})$ is the pixel intensity associated with node $i$ on the triangle element; a higher intensity value indicates greater stiffness. The finite element mesh was deformed by applying an equibiaxial displacement ($d=0.2$) to the boundaries, as shown in Fig. 1.
5.2 Hyperelastic constitutive models
We evaluated the effectiveness of our proposed PINN architecture in identifying the heterogeneous properties of soft tissues on three popular hyperelastic constitutive models: the compressible Neo-Hookean [80], incompressible Mooney Rivlin  [81], and incompressible Gent [82] models.
Neo-Hookean model. The Neo-Hookean model [80] is the simplest nonlinear elastic material model characterized by the 2D strain energy density function, $\Psi$,
where $I_{1}$ is the first principal invariants denoted as $I_{1}=\text{trace}(F^{T}\cdot F)$, $F$ is the deformation gradient denoted as $F_{ij}=\delta_{ij}+u_{i,j}$, $u_{i}$ is the displacement vector, and $\lambda$ and $\mu$ are the Lamé’s elasticity parameters. The first Piola-Kirchhoff stress of hyperelastic material models is defined as the derivative of the strain energy function with respect to the deformation gradient. In Neo-Hookean model,